• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 16:00:46 on January 03, 2008

    Entry number 222495

    keyword=Summary trigger work december 3rd 2008

    I continued working for the coincidence trigger for E06-007. Most of the time spent to find the old pieces usable of the previous coincidence trigger to be reused.

    Signals generated for coincidence trigger : -T1 output from mixed logic was sent to discriminator group 2, the output signal was set to 150 ns -T3 output from mixed logic was sent to discriminator group 4, the output signal was set to 30 ns

    the two signal were fed to the logic module to make coincidence. The T5 signal was sent to NIM/ECL module in top CAMAC crate on channel 10 ( channel 11 seems to be bad )

    coincidence trigger was fed to TS.

    Added additional delay for T1 going to TS now is going through additionnal delay RDelay channel 7 and 8 now set at 8 and 8.

    T2 was already delayed : it is on RDelay1 channel 11 to 13

    T3 was restored as previously and is on RDelay2 channel 1 and 2.

    Left to be done for the trigger : -split and delay T4 to go in TS -check retiming and gate to be compatible with T5 timing ( retiming might need to be delayed a bit to avoid forced retiming this will be done after CSR ) -check/recable coincidence time ( this will be done tomorrow during energy change )

    A copy of this log entry has been emailed to: rom@jlab.org